# Solving rescheduling problems in heterogeneous urban railway networks using hybrid quantum-classical approach

### v3

@@ 6 Computational results
In this section, we demonstrate the performance of the proposed hybrid (quantum classical) approach on a network at Polish railways - central part of the Upper Silesia Metropolis. In particular, we compare the performance of DWave’s hybrid quantum-classical solver (CQM) with CPLEX. The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different level of disturbances and disruptions including various initial delays and track closures. As the objective is tied to secondary delays, we can roughly apply it to asses the degree of difficulty of the problem.
Section 6.1 describes the considered network and traffic characteristics. Section 6.2 presents results concerning generic examples on the core line of the network, comparing double track line scenarios with single track line scenarios and closure scenarios. Section 6.3 presents analogous results for larger network with real-life timetable. Here the case studies concerns various scenarios with different degree of difficulty.
@@ 6.1 Railway network
We use the Open Railway Map [^1] https://www.openrailwaymap.org/, visited 2022-11-25 representation of the selected part of the Polish railway network located in the central part of the Metropolis GZM (Poland), presented in Fig 2. The selected part of the network combines heterogeneous network layout including single-, double-, triple-, and quadruple-track lines. It comprises 25 nodes including $11$ stations and $3$ junctions, $146$ blocks, and 2 depots. The infrastructure is managed by Polish state infrastructure manager PKP PLK (Polskie Koleje Państwowe, Polskie Linie Kolejowa eng. Polish State Railways, Polish Railway Lines).
@@
The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analysed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O). As described before, the decision stations are the only ones that appear explicitly in our model. Due to operational reasons resulting from track design, the decisions on train departure times (and thus on trains’ order) are, in the current operational practice, de facto made with respect to the decision stations. All other stations and junctions are treated as line blocks as long as no rerouting or retracking is considered; they will appear in the model implicitly via parameters. These stations include:
@@
In all computation priority weights in Eq. (2) we use $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable). As for the solver parameter t_min, after initial experiments we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality. Sensitivity analysis regarding this parameter is performed in Appendix A for selected instances.
@@ 6.2 Synthetic experiments on the selected railway line
As the first set of experiments we address, synthetic instances on a part of the network - railway line KO-GLC, see Fig 2. The goal in these scenarios is to compare the CPLEX and CQM solvers for: the double track line ($line1$), the double track line with closures ($line2$), and the single track line ($line3$). Shunting movements and rolling stock circulation are not considered in this set of experiments. In particular, the addressed configurations are the following:
@@
    $Line1$, the double track line with dense traffic. We use cyclic $3$-hour timetable with $10$ trains each hour and each direction, i.e. we have $59$ trains.
@@
    $Line3$, the whole line is considered as a single track, with a timetable of $3$ hours traffic and $21$ trains.
For $line1$ and $line3$, the original timetable is feasible, while for $line2$ it is not feasible due to the closure. For each of the lines we compute $12$ instances, each with different initial delays of trains subsets. Instance 00 yields no initial delays. Thereafter, we use the parameter value $d_{\max}=40\ \text{min}$, in all cases in this and subsequent subsection. Assuming that each train interacts with $N(d_{\max})=12$ trains for $line1$ and $line2$, and with $N(d_{\max})=6$ trains for $line3$, Eqs. (17) - (19), (22) allows us to quickly estimate the sizes of the problem. This sizes are compared with computed (comp.) actual number of variables and constraints (mean over instances) in Tab. 1. Here, $line2$ is the most complex in terms of constraints, what is reflected later by computational difficulties of some of its instances. Furthermore, $line2$ is partially double track and partially single track, hence the actual number of constraints is bounded from above by the estimate for the single track line and from below by the estimate for the double track line. $Line1$ includes double-, whereas $line3$ single-track segments only. There actual number of constraints and variables appeared to be close to the estimate.
@@
Table 1: Synthetic experiments. The first column is the ordinal number of the line. For the estimated and actual number of variables and constraints, we use $\alpha=2/3$ in Eqs. (17) - (19), (22) (doub. - double track, sing. - single track, comp. mean - computed actual number of variables or constraints, mean over instances).
@@
The computational results are presented in Figs. 3,  4, and 5 for $line1$, $line2$, and $line3$, respectively. Each figure shows objective $\times$ $d_{\text{max}}$ (top part), computation time (comp. time, middle part) and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s. CPLEX results are also displayed.
@@
For scenario $line1$ (Fig. 3), CQM solver obtained optimal solutions for all cases (except for instances $6$ and $11$). Similarly, for $line3$ (Fig. 5), only small deviations from the optimum can be observed for a number of instances. All computational times for $line1$ and $line3$ were significantly below $0.4$min, although CPLEX performed faster. Comparing CQM with t_min$=5s$ and t_min$=20s$, we have almost the same objectives but t_min$=20s$ requires almost $4$ times longer computational time.
@@
For $line2$ (Fig. 4), all CQM solutions were suboptimal, although feasible. More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than CPLEX, in several cases up to 5×5$\times$5 × faster. This difference in time is clear while setting the minimal processing time of the CQM solver t_min$=5s$ (see Section 5). This observation is approved by more detailed discussion on t_min parameter in Appendix A. In principle, $line1$ (Fig. 3) and $line3$ (Fig. 5) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster. We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems.
Figure 3: Synthetic experiments $Line1$, comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM. All the displayed solutions are feasible. Total computational time (middle panel) and QPU times (lower panel) were provided by the D-Wave output. Results of both solvers were similar in terms of the objective, but the CPLEX computation was faster.
@@
Figure 4: Synthetic experiments $Line2$, comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM. Note that the CQM is fast, even when compared to CPLEX. The found solutions are suboptimal but can be useful. All the displayed solutions are feasible. $Line2$, namely the double track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and for the quantum solver (in terms of objective). On this example, there are instances, where the current CQM D-Wave hybrid solver outperforms CPLEX in terms of computational time. Increasing the value of the t_min parameter in most cases improves the quality of the solution at the cost of computational time.
Figure 5: Synthetic experiments $Line3$, comparison of the performance of classical solvers (CPLEX) with that of the hybrid CQM. All the displayed solutions are feasible. Results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
@@ 6.3 Real-life experiments on the considered rail network
In this subsection, we will test the CQM solver on more realistic scenarios of the railway traffic on presented network. We will also focus on the track closure situation, turning a double track line segment into a single track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2. Our computations address various use cases of train delays outside the network and within the network as well as the partial track closure. We constructed $9$ networks with increased level of difficulty ranging from small delays to disruptions, i.e. it is expected more trains are involved and/or disturbances are spread more over the network. These networks are tabulated in Table  2. Their details are as follows:
@@
      Networks $4$ and $5$ concerns rerouting trains from double track line KTC-CB to the single track line with higher passing times, see Fig. 2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
@@
      Network $6$ concerns multiple closures, i.e. change of multiple line KZ-KO(STM) and double track line KO-KL to single track lines. (See Fig. 2; the reason can be e.g. upcoming reconstruction works on this part of network.)
@@
A. Disturbtions from outside the network
@@
B. Disturbtions originated within the network
@@
C. Disturbtions within the network and closures
4 11 KTC-CB 10 trains KS + IC Double track line KTC-CB closed trains rerouted to single track line
5 11 whole network most trains Double track line KTC-CB closed as in 4 + delays as in 3
6 10 whole network all trains Multiple track KZ - KO(STM) and double track KO - KL changed into single track + delays as in 3
7 11 whole network all trains Multiple track KZ - KO(STM) and double track KO - KL changed into single track + double track line KTC-CB closed as in 4 + delays as in 3
@@ C. Disturbtions within the network and closures
Table 2: Real-life experiments. Networks of particular dispatching problems. In networks 00 - $5$ most of analysed network is double tracked, from network $6$ upward the number of single track lines in the network increases at the expense of the number of double and multiple track lines. For each network we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains. Beside the #$\#S$ stations there are $2$ depots in each network.
@@
mean value over $5$ realis.
@@
Table 3: Real-life experiments. Results of ILP optimization on CPLEX and CQM D-Wave hybrid solver. $d_{\max}=40$ was set for all networks. Computational times and QPU time were reported in D-Wave’s output. We have selected t_min = $5$s as it provides the best balance between computational time and the quality of solution. For detailed analysis see Appendix A. This appendix shows a detailed analysis of the impact of t_min on computation time and solution quality for two specific real-life lines $7$ and $9$. The last columns present a percentage comparison (CQM - CPLEX ) / CPLEX, hence, the positive values represent CQM advantage, whereas negative values show CPLEX advantage.
@@
The level of difficulty increases with respect to the number of variables and constraints. Observe that CQM always returns a feasible solution. For the networks $0-3$ with initial delays only, CQM generated optimal solutions, while in the remaining networks, the objective value was somewhat higher than the actual optima obtained with CPLEX. As for computational times, for CPLEX they grow significantly - from $0.072$ to almost $8s$ (network $9$). Meanwhile those of CQM remain almost constant around $5.1s$. QPU times remain not greater than $0.032s$. Remarkably, for the hardest instances: networks $7$ and $9$, the CQM hybrid solver even outperforms CPLEX with respect to computational time. It does not find the actual optimum, just a feasible solution close to it. For example, in network $9$, the computed CQM solution is around 30% worse than the optimal one. In conclusion, we expect benefits from the application of the hybrid solver on medium scale railway networks with multiple closures, as in networks 7-9. This is in accordance with our observation on synthetic data in Fig. 4, where the benefits appeared for more complex instances, too.
To evaluate the extent of difficulty of the problems analysed in Tab. 3, Tab. 4 presents mean (over stations and realizations) delays, resulting from CQM solutions, on selected stations (i.e. stations in central part of the network in Fig. 2). Meaningful delays are introduced starting from network $4$ onward, i.e. for networks with closures in a limited number of stations, see also Tab. 2. From network $5$ onward, delays become more uniformly distributed among the stations when compared with results in Tab. 2. However, for networks $7$ $9$ the mean delays are the highest and almost uniformly distributed among all stations, showing that the spread of disruption impacts the whole network. These have appeared to be more challenging network disruptions.
Fig. 6 shows more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (top panel) vs CQM (bottom panel). For network $4$ (left panel of Fig. 6), bigger delays occur mainly at stations KO(STM), KO and CB for both CPLEX and CQM solutions, albeit for the former are somewhat smaller. This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Tab. 2), and thus the spread is limited to these 3 stations. For network $7$ (right panel of Fig. 6), delays are more spread through the network due to disruption and initial disturbances. To conclude, we can expect advantage from a hybrid (quantum classical) approach particularly in case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network.
@@
Table 4: Real-life experiments. Mean (over trains and realizations) secondary delay resulted from CQM solutions (Tab. 3) for all networks in Tab. 2. Non-zero values for network 00 comes from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes delay but make it up and leaves the network on time - these are typical issues for the railway traffic on analysed network.
@@
Figure 6: Real-life experiments. Statistics of CPLEX (top panel) and CQM (bottom panel) solutions (Tab. 3) for selected networks in Tab. 2. The statistics were calculated over trains in the classical cases, and over trains and realizations in the CQM cases. Vertical lines are ranges, while horizontal lines are $0.25$ and $0.75$ percentiles.

### v4

@@ 6 Computational results
In this section, we demonstrate the performance of the proposed hybrid (quantum-classical) approach on a network at Polish railways - the central part of the Upper Silesia Metropolis. In particular, we compare the performance of D-Wave’s hybrid quantum-classical solver (CQM) with CPLEX. The comparison is performed for diverse network layouts including single-, double-, and multiple-track lines, and different level of disturbances including various initial delays and track closures. As the objective is tied to secondary delays, we can roughly apply it to assess the degree of difficulty of the problem.
Section 6.1 describes the considered network and traffic characteristics. Section 6.2 presents results concerning generic examples on the core line of the network, comparing double-track-line scenarios with single-track-line scenarios and closure scenarios. Section 6.3 presents analogous results for larger network with real-life timetable. Here the case studies concern various scenarios with different degrees of difficulty.
@@ 6.1 Railway network
We use the Open Railway Map [^1] https://www.openrailwaymap.org/, visited 2022-11-25 representation of the selected part of the Polish railway network located in the central part of the Metropolis GZM (Poland), presented in Fig 2. The selected part of the network combines a heterogeneous network layout including single-, double-, triple-, and quadruple-track lines. It comprises 25 nodes including $11$ stations and $3$ junctions, $146$ blocks, and 2 depots. The infrastructure is managed by Polish state infrastructure manager PKP PLK (Polskie Koleje Państwowe, Polskie Linie Kolejowa eng. Polish State Railways, Polish Railway Lines).
@@
The decision stations include stations {KO, KO(STM), CB, KL}, stations bounding the analyzed part of the network {GLC, CM, KZ, Ty, Mi}, depots {KO(KS), KO(IC)} from which regional and intercity trains are being shunted to and from {KO}, and stations on the single-track line – {MJ} to allow meet and pass (M-P) or meet and overtake (M-O). As described before, the decision stations are the only ones that appear explicitly in our model. Due to operational reasons resulting from track design, the decisions on train departure times (and thus on trains’ order) are, in the current operational practice, de facto made with respect to the decision stations. All other stations and junctions are treated as line blocks as long as no rerouting or retracking is considered; they will appear in the model implicitly via parameters. These stations include:
@@
In all computation priority weights in Eq. (2) we use $w_{j}=1$ for stopping (local) trains $w_{j}=1.5$ for intercity (fast) trains $w_{j}=1.75$ for express trains, and $w_{j}=0$ for shunting (if applicable). As for the solver parameter t_min, after initial experiments, we have decided to prefer t_min$=5$s as it provides a good balance between computation time and expected solution quality. A sensitivity analysis regarding this parameter will be performed in Section 6.3.
@@ 6.2 Synthetic experiments on the selected railway line
As the first set of experiments we address, synthetic instances on a part of the network - railway line KO-GLC, see Fig 2. The goal in these scenarios is to compare the CPLEX and CQM solvers for: the double-track line ($line1$), the double-track line with closures ($line2$), and the single-track line ($line3$). Shunting movements and rolling stock circulation are not considered in this set of experiments. In particular, the addressed configurations are the following:
@@
    $Line1$, the double-track line with dense traffic. We use cyclic $3$-hour timetable with $10$ trains each hour and each direction, i.e. we have $59$ trains.
@@
    $Line3$, the whole line is considered as a single-track, with a timetable of $3$ hours traffic and $21$ trains.
For $line1$ and $line3$, the original timetable is feasible, while for $line2$ it is not feasible due to the closure. For each of the lines, we compute $12$ instances, each with different initial delays of train subsets. Instance 00 yields no initial delays. Thereafter, we use the parameter value $d_{\max}=40\ \text{min}$, in all cases in this and subsequent subsection. Assuming that each train interacts with $N(d_{\max})=12$ trains for $line1$ and $line2$, and with $N(d_{\max})=6$ trains for $line3$, Eqs. (17) - (19), (22) allows us to quickly estimate the sizes of the problem. These sizes are compared with the computed (comp.) actual number of variables and constraints (mean over instances) in Tab. 1. Here, $line2$ is the most complex in terms of constraints, which is reflected later by the computational difficulties of some of its instances. Furthermore, $line2$ is partially double-track and partially single-track, hence the actual number of constraints is bounded from above by the estimate for the single-track line and from below by the estimate for the double-track line. $Line1$ includes double-, whereas $line3$ single-track segments only. There actual number of constraints and variables appeared to be close to the estimate.
@@
Table 1: Synthetic experiments. The first column is the ordinal number of the line. For the estimated and actual number of variables and constraints, we use $\alpha=2/3$ in Eqs. (17) - (19), (22) (doub. - double-track, sing. - single-track, comp. mean - computed actual number of variables or constraints, mean over instances).
@@
The computational results are presented in Fig. 3 a) b) c) for $line1$, $line2$, and $line3$, respectively. Each figure shows objective $\times$ $d_{\text{max}}$ (top part), computation time (comp. time, middle part), and QPU access time (bottom part) for CQM hybrid solver with t_min equal $5$ and $20$s. CPLEX results are also displayed.
@@
For scenario $line1$ (Fig. 3 a) ), the CQM solver obtained optimal solutions for all cases (except for instances $6$ and $11$). Similarly, for $line3$ (Fig. 3 c) ), only small deviations from the optimum can be observed for a number of instances. All computational times for $line1$ and $line3$ were significantly below $0.4$min, although CPLEX performed faster. Comparing CQM with t_min$=5s$ and t_min$=20s$, we have almost the same objectives but t_min$=20s$ requires almost $4$ times longer computational time.
@@
For $line2$ (Fig. 3 b) ), all CQM solutions were suboptimal, although feasible. More remarkably, for $9$ instances at t_min$=5s$ and $3$ instances at t_min$=20s$ (out of $12$ instances) CQM was faster than CPLEX, in several cases up to 5×5$\times$5 × faster. This difference in time is clear while setting the minimal processing time of the CQM solver t_min$=5s$ (see Section 5). In principle, $line1$ (Fig. 3 a) ) and $line3$ (Fig. 3 c) ) are less complicated in terms of number of constraints, and here solutions of CPLEX and CQM are similar in terms of objective, but CPLEX is faster. We can conclude that the utility of the CQM solver can be expected for the more complex instances of railway scheduling problems.
@@
Figure 3: Synthetic experiments $Line1$, $Line2$, $Line3$ comparison of the performance of classical solver (CPLEX) with that of the hybrid CQM. All the displayed solutions are feasible. Total computational time (middle panel) and QPU times (lower panel) were provided by the D-Wave output. Panel a) - $line1$, results of both solvers were similar in terms of the objective, but the CPLEX computation was faster. Panel b) - $line2$ namely the double-track line with closures and dense traffic, has appeared to be most challenging for both the classical solvers (in terms of computational time) and for the quantum solver (in terms of objective). On this example, there are instances, where the current CQM D-Wave hybrid solver outperforms CPLEX in terms of computational time. Increasing the value of the t_min parameter in most cases improves the quality of the solution at the cost of computational time. Panel c) - $line3$ all the displayed solutions are feasible. Results of both solvers were similar in terms of the objective, but CPLEX computation was faster.
@@ 6.3 Real-life experiments on the considered rail network
In this subsection, we will test the CQM solver on more realistic scenarios of the railway traffic on the presented network. We will also focus on the track closure situation, turning a double-track line segment into a single-track one, under dense traffic, to elaborate our findings from $line2$ in Section 6.2. Our computations address various use cases of train delays outside the network and within the network as well as the partial track closure. We constructed $9$ networks with increased levels of difficulty ranging from small delays to disruptions, i.e. it is expected more trains are involved and/or disturbances are spread more over the network. These networks are tabulated in Table  2. Their details are as follows:
@@
      Networks $4$ and $5$ concerns rerouting trains from double-track line KTC-CB to the single-track line with higher passing times, see Fig. 2 (trains have no initial delays in network $3$ and some initial delays in network $4$).
@@
      Network $6$ concerns multiple closures, i.e. change of multiple line KZ-KO(STM) and double-track line KO-KL to single-track lines. (See Fig. 2; the reason can be e.g. upcoming reconstruction works on this part of network.)
@@
A. Disturbances from outside the network
@@
B. Disturbances originated within the network
@@
C. Disturbances within the network and closures
4 11 KTC-CB 10 trains KS + IC double-track line KTC-CB closed trains rerouted to single-track line
5 11 whole network most trains double-track line KTC-CB closed as in 4 + delays as in 3
6 10 whole network all trains multiple-track KZ - KO(STM) and double-track KO - KL changed into single-track + delays as in 3
7 11 whole network all trains multiple-track KZ - KO(STM) and double-track KO - KL changed into single-track + double-track line KTC-CB closed as in 4 + delays as in 3
@@ C. Disturbances within the network and closures
Table 2: Real-life experiments. Networks of particular dispatching problems. In networks 00 - $5$ most of analyzed network is double-tracked, from network $6$ upward the number of single-track lines in the network increases at the expense of the number of double and multiple-track lines. For each network we considered a two-hour time with an afternoon rush-hour timetable from the year $2021$, with #$\#J=27$ trains. Beside the #$\#S$ stations there are $2$ depots in each network.
@@
mean value over $5$ realiz.
@@
Table 3: Real-life experiments. Results of ILP optimization on CPLEX and CQM D-Wave hybrid solver. $d_{\max}=40$ was set for all networks. Computational times and QPU time were reported in D-Wave’s output. We have selected t_min = $5$s as it provides the best balance between computational time and the quality of solution. The last columns present a percentage comparison (CQM - CPLEX ) / CPLEX, hence, the positive values represent CQM advantage, whereas negative values show CPLEX advantage.
@@
The level of difficulty increases with the number of variables and constraints. Observe that CQM always returns a feasible solution. For the networks $0-3$ with initial delays only, CQM generated optimal solutions, while in the remaining networks, the objective value was somewhat higher than the actual optima obtained with CPLEX. As for computational times, for CPLEX they grow significantly - from $0.072$ to almost $8s$ (network $9$). Meanwhile, those of CQM remain almost constant around $5.1s$. QPU times remain not greater than $0.032s$. Remarkably, for the hardest instances: networks $7$ and $9$, the CQM hybrid solver even outperforms CPLEX with respect to computational time. It does not find the actual optimum, just a feasible solution close to it. For example, in network $9$, the computed CQM solution is around 30% worse than the optimal one. In conclusion, we expect benefits from the application of the hybrid solver on medium-scale railway networks with multiple closures, as in networks 7-9. This finding follows our observation on synthetic data in Fig. 3 b), where the benefits appeared for more complex instances, too.
To evaluate the extent of difficulty of the problems analyzed in Tab. 3, Tab. 4 presents mean (over stations and realizations) delays, resulting from CQM solutions, on selected stations (i.e. stations in the central part of the network in Fig. 2). Meaningful delays are introduced starting from network $4$ onward, i.e. for networks with closures in a limited number of stations, see also Tab. 2. From network $5$ onward, delays become more uniformly distributed among the stations when compared with results in Tab. 2. However, for networks $7$ $9$ the mean delays are the highest and almost uniformly distributed among all stations, showing that the spread of disruption impacts the whole network. These have appeared to be more challenging network disruptions.
Fig. 4 shows a more detailed statistical analysis for two networks $4$ and $7$ and compares CPLEX (left, red) vs CQM (right, blue). For network $4$, bigger delays occur mainly at stations KO(STM), KO and CB for both CPLEX and CQM solutions, albeit for the former are somewhat smaller. This reflects the initial disturbances, which indeed occurred mainly between KO and CB (see Tab. 2), and thus the spread is limited to these 3 stations. For network $7$, delays are more spread through the network due to disruption and initial disturbances. To conclude, we can expect an advantage from a hybrid (quantum-classical) approach, particularly in case of more complex problems with extensive disruptions and disturbances occurring throughout the whole network. To facilitate further evaluation of the solutions, train diagrams are presented in Appendix A.
@@
Table 4: Real-life experiments. Mean (over trains and realizations) secondary delay resulted from CQM solutions (Tab. 3) for all networks in Tab. 2. Non-zero values for network 00 come from counted delays in shunting movement in KO and KO(STM) that are not included in the objective, as well as some cases where the train gets a few minutes, delay but makes it up and leaves the network on time - these are typical issues for the railway traffic on analyzed network.
@@
Figure 4: Real-life experiments. Statistics of CPLEX (left, read) and CQM (right, blue) solutions (Tab. 3) for selected networks in Tab. 2. The statistics were calculated over trains in the classical cases, and over trains and realizations in the CQM cases. Vertical lines are ranges, while horizontal lines are $0.25$ and $0.75$ percentiles.
The right choice of the t_min parameter can improve results meaningfully. Let us analyze in more detail the role of the t_min solver parameter for selected problems. In particular, for networks $7$ and $9$ therein, we have performed runs with t_min sweeping over a range $[5,60]$; the results are presented in Fig. 5 a) and b). The objective value is roughly log-linear in computational time. For small t_min parameter values the computational time is shorter, whereas for large t_min it grows exponentially. Overall, the improvement in the objective is overwhelmed by the increase of computational time. This is important from the point of view of the algorithm layout, as it may be tied to a power law scaling. Such behavior is plausible in the case of QA methods (Soriani et al., 2022; Albash et al., 2017; Domino et al., 2022a).
Figure 5: Real-life experiments. The sweep over t_min parameter for network $7$ a) and network $9$ b), see Tab. 2. All solutions were feasible. For higher t_min we expect smaller objective values but higher total computational time. For each parameter value, there were $5$ realizations of the experiment were performed. Interestingly, a linear (negatively sloped) relation between the objective and the logarithm of ${\tt t\_min}$ can be observed, suggesting power law scaling.
7 Conclusions
We have demonstrated that quantum annealers can be readily applicable in train rescheduling optimization in urban railway networks. In particular, we have encoded rescheduling problems in a linear integer program and solved these using a state-of-the-art solver (CPLEX), as well as the QA based CQM hybrid (quantum-classical) solver. The results suggest that quantum computing and QA in particular, although an early-stage technology, are ready to tackle challenging railway rescheduling problems. While the CQM hybrid solver does not outperform classical solvers in general, we have found specific cases in which they were actually better. This supports the expectation that future quantum devices will be efficient in solving large-scale problems, and in our case, real-time railway dispatching problems on the national scale; even in the range that is beyond the scope of current exact models and heuristics.
While the quantum-based solvers possibly return suboptimal solutions, these are still feasible and often better than the solutions obtained manually or based on smaller-scale models. In each computation, the CQM hybrid solver always has taken some advantage of the QPU. Hence, such a little help from the QPU must have boosted the classical heuristics in the hybrid solver, even if quantum annealer devices are subjects of many ongoing debates.
Several future research directions can be determined. The first concerns the ongoing worldwide works on the development of fault-tolerant quantum computing devices (DiVincenzo, 2000) with proper error correction/mitigation, this will boost results of approaches supported by quantum computing (and QA in particular). The second concerns the creation of a custom open-source hybrid solver that is dedicated to railway problems and will use the noise of quantum device to our advantage e.g. by modeling some stochastic behavior on railways. The third concerns the evolution and application of quantum-inspired techniques such as tensor networks technique, and simulated bifurcation (Goto, 2019). With all these future developments, QA-based approaches tend to become more powerful and ready to address for the first time large-scale real-time challenges in railway traffic management, which are currently unsolvable with traditional techniques.
Data availability
Acknowledgments
This research was supported by the Ministry of Culture and Innovation and the National Research, Development and Innovation Office within the Quantum Information National Laboratory of Hungary (Grant No. 2022-2.1.1-NL-2022-00004) (MK), and by the Silesian University of Technology Rector’s Grant no. BKM - 720/RT2/2023 12/020/BKM_2023/0252 (KK). This research was funded in part by National Science Center, Poland, under grant number 2019/33/B/ST6/0201 (LB) and 2023/07/X/ST6/00396 (KD). For the purpose of Open Access, the authors has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
We would like to thank Sebastian Deffner as well as the whole scientific community of the Department of Physics, University of Maryland, Baltimore County (UMBC) for valuable discussion; and to Bartłomiej Gardas, and Zbigniew Puchała for valuable motivation; and Katarzyna Gawlak, Akash Kundu, and Özlem Salehi for data validation and assistance with coding. We acknowledge the cooperation with Koleje Ślaskie sp. z o.o. (eng. Silesian Railways).
Appendix A. Train diagrams of CPLEX and CQM